OpenStudy (anonymous):

int e^2x(cos3x)dx please help :)

OpenStudy (anonymous):

:( lol

OpenStudy (anonymous):

u=cos3x?

OpenStudy (anonymous):

but i feel like there might be some trig identity at play here

OpenStudy (anonymous):

ah yeah batman

OpenStudy (anonymous):

At first glance integration by parts comes to mind.

OpenStudy (anonymous):

ok so if i set u=cos3x

OpenStudy (anonymous):

Have you tried? Let u = cos3x

OpenStudy (anonymous):

then du is what 1/3(sin3x)?

OpenStudy (xapproachesinfinity):

integration by part is what i think of when i see this

OpenStudy (anonymous):

You're taking the derivative

OpenStudy (anonymous):

Not integrating cos3x

OpenStudy (anonymous):

so 3sin3x?

OpenStudy (anonymous):

Try again :)

OpenStudy (anonymous):

sin(3x)*(3x)' right?

OpenStudy (anonymous):

or is it -sin

OpenStudy (anonymous):

Hehe nice

OpenStudy (anonymous):

lmao GRR TRIGGGG

OpenStudy (anonymous):

v = ?

OpenStudy (anonymous):

v=e2x?

OpenStudy (xapproachesinfinity):

good going^_^

OpenStudy (anonymous):

Ye homeslice dawg my mang dis is how we do 1t

OpenStudy (anonymous):

omg batman just got real that's what im talking bout

OpenStudy (jhannybean):

think of... LIPETs. (log, integral, polynomial, exponential, trig) when doing partials (delayed post)

OpenStudy (xapproachesinfinity):

LIPETs hit me back with some memories^_^!!!

OpenStudy (jhannybean):

:)

OpenStudy (anonymous):

ok wow so i end up with \[e ^{2x}\cos(3x)-\int\limits_{}^{}-3e^{2x}(sinx)dx\]

OpenStudy (anonymous):

so then i take out the -3

OpenStudy (xapproachesinfinity):

sin3x not sinx

OpenStudy (anonymous):

oh yeah sorry. thats what i have on paper, just made the eq bad over here

OpenStudy (xapproachesinfinity):

this a hick of a intergral by part you will repeat the integration several times

OpenStudy (anonymous):

so i have now

OpenStudy (anonymous):

\[e^{2x}\cos(3x)+3e^{2x}\sin(3x)-9\int\limits_{}^{}e^{2x}\cos(3x)\]

OpenStudy (xapproachesinfinity):

Usually this problem you continue till something repeats and thereafter you need to cancel some term to finish it off

OpenStudy (anonymous):

can i just add the 9inte^2xcos3x to the other side

OpenStudy (xapproachesinfinity):

very good observation my friend! that's how it work you throw that term to the other side

OpenStudy (anonymous):

sweet. just throwing me off cause that 9 lol.

OpenStudy (xapproachesinfinity):

I'm not sure you did everything correctly! but you steps are descent and that's how to throw off this intergration

OpenStudy (xapproachesinfinity):

this one turned to be less steps than i anticipated! As I said some of them you need to carry on tell you recognize repeatition

OpenStudy (jhannybean):

Then you just label the repetition with a letter.

OpenStudy (anonymous):

label the repetition with a letter?

OpenStudy (xapproachesinfinity):

Yeah like a big Letter A say A=........-9A A is your starting integral or could be something else

OpenStudy (xapproachesinfinity):

you are doing great you know! continue....

OpenStudy (anonymous):

i ended with (1/10)(e^2x)(cos3x+3sin3x)+c but wolfram says it should be (1/13)(e^2x)(2cos3x+3sin3x)+c

OpenStudy (anonymous):

and i can't figure out what i did wrong

OpenStudy (anonymous):

i started with ine^2xcos3x u=cos3x du=-3sin3x dv=e^2xdx v=e^2x

OpenStudy (anonymous):

so then i got e^2xcos3x+3inte^2x(sin3x)dx u=sin3x du=3cos3x dv=e^2xdx v=e^2x

OpenStudy (anonymous):

so e^2xcos3x+3[e^2xsin3x-int3e^2xcosxdx]

OpenStudy (xapproachesinfinity):

just go back to your calculations and see if there is something wrong and beside i don't really trust wolfram always but for this one i this wolfram is right the problem is that wolfram goes deep in integrating